# Incubulable hyperbolic $3$-pseudomanifold groups

### v1

@@ 5. Proofs of the main results
We may therefore apply Theorem 2.2 ([KM, Theorem A]) to obtain a negatively curved metric $\hat{d}$ on $\widehat{M}$. This proves (1).
@@
so the set $H^{\prime}$ satisfies all the conditions of Theorem 2.4 ([KM, Theorem B]). In particular, $\widehat{H^{\prime}}$ is isotopic to a locally convex, and hence $\pi_{1}$-injective, subset of $\widehat{M}$. But $\pi_{1}(\widehat{H^{\prime}})\cong\pi_{1}(\widehat{T}_{n})$ has property (T) by Lemma 2.8. This establishes item (2) of the theorem.
Finally, to prove (3), note that the systole of $\partial M$ goes to infinity with $n$. On the other hand, the systole of a closed hyperbolic surface is bounded by a function of the genus (see [MU71]), so the genus of the boundary components must be going to infinity as well. Moreover, the Euler characteristic of $\widehat{M}_{n}$ is the sum of the genera of the links of the singular points. ∎
The statements in Corollary 1.4 follow from the existence of an infinite subgroup with property (T) via standard arguments. We include a proof for the reader’s convenience. Note that in our context property (T) is equivalent to property (FH) by the Delorme–Guichardet Theorem, see [BHV08, Theorem 2.12.4].
@@
Let $K$ be the infinite subgroup of $\pi_{1}(\widehat{M}_{n})$ with property (T). A group with property (T) has a global fixed point whenever it acts cubically on a (not necessarily finite dimensional) $\textrm{CAT}(0)$ cube complex. (Special cases of this fact were proved first by Niblo–Reeves [NR97] and Niblo–Roller [NR98]; the statement in full generality is proved by Cornulier in [CO13].) In particular if $\pi_{1}(\widehat{M}_{n})$ acts on a $\textrm{CAT}(0)$ cube complex, then $K$ must fix a point, and so the action is not proper. This establishes conclusion (1).
Similarly, since $K$ has property (FH), [BHV08, Corollary 2.7.3] implies that any action of $K$ by isometries on a real or complex hyperbolic space has a global fixed point, which proves (2).
Finally, to prove (3) assume by contradiction that $\pi_{1}(\widehat{M}_{n})$ has a RFRS subgroup of finite index $G^{\prime}$. Then $K^{\prime}=K\cap G^{\prime}$ is a finite index subgroup of $K$, hence inherits property (T); see [BHV08]. Moreover, $K^{\prime}$ is RFRS because it is a subgroup of $G^{\prime}$, so $K^{\prime}$ surjects $\mathbb{Z}$; see [AG08]. But $K^{\prime}$ has property (T), so its abelianization is finite (see [BHV08, Corollary 1.3.6]), which leads to a contradiction. ∎
Finally, let us prove the claim in Remark 1.7
@@ 6. Questions
We showed in Corollary 1.4 that our $3$-pseudomanifold groups are not virtually RFRS. We ask if they are residually finite. (It is a well-known question whether every word hyperbolic group is residually finite.) Note that the groups with property (T) we use in this paper (see §2.3) are finite index subgroups of some of the groups constructed in [LMW19], and even those are not known to be residually finite.
@@
Finally, we ask whether cubulability can be ensured if the metric is sufficiently nice, say polyhedral, i.e., piecewise hyperbolic. Note that the metric we use from [KM] is not polyhedral.
@@
By Remark 1.7 the groups constructed in this paper do not have property (T), and by [FU99] an infinite $3$-manifold group never has property (T).

### v2

@@ 5. Proofs of the main results
We may therefore apply Theorem 2.2 ([33, Theorem A]) to obtain a negatively curved metric $\hat{d}$ on $\widehat{M}$. This proves (1).
@@
so the set $H^{\prime}$ satisfies all the conditions of Theorem 2.4 ([33, Theorem B]). In particular, $\widehat{H^{\prime}}$ is isotopic to a locally convex, and hence $\pi_{1}$-injective, subset of $\widehat{M}$. But $\pi_{1}(\widehat{H^{\prime}})\cong\pi_{1}(\widehat{T}_{n})$ has property (T) by Lemma 2.8. This establishes item (2) of the theorem.
Finally, to prove (3), note that the systole of $\partial M$ goes to infinity with $n$. On the other hand, the systole of a closed hyperbolic surface is bounded by a function of the genus (see [42]), so the genus of the boundary components must be going to infinity as well. Moreover, the Euler characteristic of $\widehat{M}_{n}$ is the sum of the genera of the links of the singular points. ∎
The statements in Corollary 1.4 follow from the existence of an infinite subgroup with property (T) via standard arguments. We include a proof for the reader’s convenience. Note that in our context property (T) is equivalent to property (FH) by the Delorme–Guichardet Theorem, see [4, Theorem 2.12.4].
@@
Let $K$ be the infinite subgroup of $\pi_{1}(\widehat{M}_{n})$ with property (T). A group with property (T) has a global fixed point whenever it acts cubically on a (not necessarily finite dimensional) $\textrm{CAT}(0)$ cube complex. (Special cases of this fact were proved first by Niblo–Reeves [44] and Niblo–Roller [43]; the statement in full generality is proved by Cornulier in [14].) In particular if $\pi_{1}(\widehat{M}_{n})$ acts on a $\textrm{CAT}(0)$ cube complex, then $K$ must fix a point, and so the action is not proper. This establishes conclusion (1).
Similarly, since $K$ has property (FH), [4, Corollary 2.7.3] implies that any action of $K$ by isometries on a real or complex hyperbolic space has a global fixed point, which proves (2).
Finally, to prove (3) assume by contradiction that $\pi_{1}(\widehat{M}_{n})$ has a RFRS subgroup of finite index $G^{\prime}$. Then $K^{\prime}=K\cap G^{\prime}$ is a finite index subgroup of $K$, hence inherits property (T); see [4]. Moreover, $K^{\prime}$ is RFRS because it is a subgroup of $G^{\prime}$, so $K^{\prime}$ surjects $\mathbb{Z}$; see [1]. But $K^{\prime}$ has property (T), so its abelianization is finite (see [4, Corollary 1.3.6]), which leads to a contradiction. ∎
Next we prove the claim in Remark 1.7.
@@
Finally, we prove the claim in Remark 1.8. To state and prove it, we must use some notions from the theory of relatively hyperbolic groups [26, 22, 8]. A quick introduction to most of the language as we use it can be found in [39, Section 2.2]. See [20] for the definition of a relatively geometric action. If $B$ is the mirrored polyhedron from Subsection 3.1, we note that the pair $(\pi_{1}^{\operatorname{orb}}\left(B\right),\{\pi_{1}^{\operatorname{orb}}\left(\partial B\right)\})$ is relatively hyperbolic, with Bowditch boundary homeomorphic to a 22–sphere. Any finite index subgroup $\Gamma$ of $\pi_{1}^{\operatorname{orb}}\left(B\right)$ (for example the fundamental group of one of our manifolds $M_{n}$) receives an induced peripheral structure $\mathcal{P}$ so that $(\Gamma,\mathcal{P})$ is relatively hyperbolic and so the inclusion $\Gamma\to\pi_{1}^{\operatorname{orb}}\left(B\right)$ induces an identification of Bowditch boundaries. The peripheral groups $P\in\mathcal{P}$ are all commensurable to $\pi_{1}^{\operatorname{orb}}\left(\partial B\right)$. We write $(\Gamma,\mathcal{P})<(\pi_{1}^{\operatorname{orb}}\left(B\right),\{\pi_{1}^{\operatorname{orb}}\left(\partial B\right)\})$ for this relationship.
Proposition 5.2.
Let $(\Gamma,\mathcal{P})<(\pi_{1}^{\operatorname{orb}}\left(B\right),\{\pi_{1}^{\operatorname{orb}}\left(\partial B\right)\})$ be finite index. Let $X$ be a $\textrm{CAT}(0)$ cube complex, and let $(\Gamma,\mathcal{P})\curvearrowright X$ be a relatively geometric action. Then there exist $P\in\mathcal{P}$ and a hyperplane stabilizer $H$ such that $H\cap P$ is infinite.
Proof.
By contradiction, suppose there is a relatively geometric action by $\Gamma$ on a cube complex $X$ so that hyperplane stabilizers intersect parabolic subgroups only in finite subgroups.
The hyperplane stabilizers suffice to separate points at infinity of the Bowditch boundary of $(\Gamma,\mathcal{P})$ by [19, Theorem 1.4]. Their images under the inclusion of $\Gamma\to\pi_{1}^{\operatorname{orb}}\left(B\right)$ therefore separate points at infinity of the Bowditch boundary of $(\pi_{1}^{\operatorname{orb}}\left(B\right),\{\pi_{1}^{\operatorname{orb}}\left(\partial B\right)\})$, which we have noted is the same as that of $(\Gamma,\mathcal{P})$. The result [21, Theorem 1.3] then produces a relatively geometric action $\pi_{1}^{\operatorname{orb}}\left(B\right)\curvearrowright Y$ with hyperplane stabilizers commensurable to those of $\Gamma\curvearrowright X$. (Note that here we need not refine the peripheral structure, since $\pi_{1}^{\operatorname{orb}}\left(\partial B\right)$ is one-ended.) Because the hyperplane stabilizers in $\pi_{1}^{\operatorname{orb}}\left(B\right)\curvearrowright Y$ are commensurable to the hyperplane stabilizers in $\Gamma\curvearrowright X$, they have finite intersection with conjugates of $\pi_{1}^{\operatorname{orb}}\left(\partial B\right)$.
Let $v$ be a vertex of $Y$ with infinite stabilizer, and let $e$ be an edge incident to $v$. Since $\operatorname{Stab}(e)$ is (up to index $2$) equal to the intersection of $\operatorname{Stab}(v)$ with the stabilizer of the hyperplane dual to $e$, we deduce $\operatorname{Stab}(e)$ is finite. In particular the action of $\operatorname{Stab}(v)$ on $\operatorname{lk}\left(v,Y\right)$ is proper. Since the action is relatively geometric, it is cocompact, so $\operatorname{Stab}(v)$ also acts cocompactly on $\operatorname{lk}\left(v,Y\right)$. It follows that there are only finitely many conjugacy classes of elements of $\operatorname{Stab}(v)$ which move a vertex of $\operatorname{lk}\left(v,Y\right)$ a combinatorial distance $3$ or less. Any element of $\operatorname{Stab}(v)$ is an element of $\pi_{1}^{\operatorname{orb}}\left(B\right)$ so it acts on $\mathbb{H}^{3}$, with some translation length. Let $u$ be the maximum of the (finitely many) translation lengths in $\mathbb{H}^{3}$ of elements of $\pi_{1}^{\operatorname{orb}}\left(B\right)$ which stabilize some vertex $v$ and move some vertex of $\operatorname{lk}\left(v,Y\right)$ a distance $3$ or less.
We fix $n$ large enough so that Theorem 1.3 applies and so that moreover the systole of every boundary component of $M_{n}$ is strictly larger than $u$. We can restrict the action $\pi_{1}^{\operatorname{orb}}\left(B\right)\curvearrowright Y$ to the finite index subgroup $\pi_{1}(M_{n})$. The restricted action is still relatively geometric (with respect to the induced peripheral structure), and in particular cocompact. By the choice of $n$, the following condition on the action $\pi_{1}(M_{n})\curvearrowright Y$ holds:
($\dagger$) For every vertex $v$ of $Y$, every element of $\operatorname{Stab}(v)$ acts on $\operatorname{lk}\left(v,Y\right)$ by moving every vertex a combinatorial distance of at least 4.
[…]
To see that $Z$ is a $\textrm{CAT}(0)$ cube complex, we must check that $Z$ is a simply connected cube complex and satisfies Gromov’s link condition (that the link of every vertex is flag). That $Z$ is a simply connected cube complex follows from [27, Theorem 4.1 and Proposition 4.3]. Let $\overline{v}$ be a vertex of $Z$ which is the image of a vertex $v$ in $Y$. Since $Y$ is $\textrm{CAT}(0)$, $\operatorname{lk}\left(v,X\right)$ is flag. The stabilizer of $v$ in $\pi_{1}(M_{n})$ acts freely on this link, and by ($\dagger$5) it moves every vertex a combinatorial distance of at least $4$. Thus the quotient $\operatorname{lk}\left(\overline{v},Z\right)$ is also flag. ∎
@@ 6. Questions
We showed in Corollary 1.4 that our $3$-pseudomanifold groups are not virtually RFRS. We ask if they are residually finite. (It is a well-known question whether every word hyperbolic group is residually finite.) Note that the groups with property (T) we use in this paper (see §2.3) are finite index subgroups of some of the groups constructed in [37], and even those are not known to be residually finite.
@@
Finally, we ask whether cubulability can be ensured if the metric is sufficiently nice, say polyhedral, i.e., piecewise hyperbolic. Note that the metric we use from [33] is not polyhedral.
@@
By Remark 1.7 the groups constructed in this paper do not have property (T), and by [25] an infinite $3$-manifold group never has property (T).